Jelly beans and mints, worth $2.10/kg and $2.70/kg respectively, were mixed to make 500kg of mixture, which sold for $2.52/kg. How many kilograms of jelly beans were sold?
Answer by mananth(16946)

Price ---------------- quantity
jelly beans $2.70 ---------------- x kgs
mints $2.10%------ 500 - x kgs
Mixture 2.52% ---------------- 500
Total 500 kgs
2.7 x + 2.1 ( 500 - x ) = 1260
2.7 x + 1050 - 2.1 x = 1260
2.7 x - 2.1 x = 1260 - -1050
0.6 x = 210
/ 0.6
x = 350 kgs $2.70 jelly beans
150 kgs $2.10 mints

m.ananth@hotmail.ca